The length of a rectangle is 3 feet more than twice its width, and 2 feet less than a diagonal's length. Find the length of the diagonal in simplest radical form.
Answer by stanbon(48535)

The length of a rectangle is 3 feet more than twice its width, and 2 feet less than a diagonal's length. Find the length of the diagonal in simplest radical form.
---------------
Let width = W.
Then Length = "2W+3"
---------------------------
Equation:
diagonal^2 = leg^2 + leg^2
d^2 = x^2 + (2x+3)^2
d^2 = x^2 + 4x^2 + 12x + 9
d^2 = 5x^2 + 12x + 9
-----
d = sqrt(5x^2+12x+9)
===========================
Cheers,
Stan H.